# Braid positive surgery diagrams

### 27 Mar 2026 v2

@@ Abstract.
In this short note, we prove that every closed, oriented, connected $3$-manifold arises as Dehn surgery along a braid positive link.
@@ 1. Introduction
A classical result by Lickorish and Wallace [LIC62, WAL61] states that every closed, oriented, connected $3$-manifold $M$ arises as Dehn surgery along a link $L$ in $S^{3}$. That is, $M$ can be obtained by removing an open tubular neighbourhood $\nu L$ of $L$ from $S^{3}$ and gluing back to each torus boundary component of $S^{3}\setminus\nu L$ a solid torus by an orientation-preserving diffeomorphism of its boundary. Our motivation in this note stems from the following type of question.
@@
Let $\mathcal{L}$ be a set of links in $S^{3}$. Can each closed, oriented, connected $3$-manifold be obtained by Dehn surgery along a link $L\in\mathcal{L}$?
Considering the first homology of a $3$-manifold obtained by Dehn surgery directly implies that the number of components of links in $\mathcal{L}$ is unbounded if $\mathcal{L}$ gives a positive answer to Question 1. Question 1 was answered affirmatively for the following sets of links:
@@
In contrast to these results, not every closed, oriented, connected $3$-manifold can be obtained by Dehn surgery along a torus link since Dehn surgeries along torus links are never hyperbolic [MOS71, KS10]. Moreover, the answer to Question 1 is negative for the set of Montesinos links by [IJ09] and thus also for pretzel links. Question 1 also has a negative answer for 2-bridge links since these have at most two components.
@@
Every closed, oriented, connected $3$-manifold $M$ can be obtained by Dehn surgery along a braid positive link $L$ in $S^{3}$. Moreover, $L$ can be chosen to have at most $s(M)+1$ many components.
@@
All $3$-manifolds are assumed to be smooth and oriented, and are considered up to orientation-preserving diffeomorphism. For us, a link in a $3$-manifold $M$ is a smooth, closed $1$-dimensional submanifold of $M$. Surgery coefficients (see Section 2) are measured with respect to the Seifert framing.
@@ Acknowledgements
This project started at the conference on Trisections and related topics, October 2025, CIRM, Marseille, and was finished at the workshop on Knotted surfaces in $4$-manifolds, December 2025, University Regensburg. We thank Stefan Friedl for helpful comments on a first draft.
@@ 2. Surgery diagrams and Rolfsen twists
First, let us briefly review the basics of Dehn surgery. For more details, the reader is referred, for example, to [ROL03]. Let $L=L_{1}\cup\dots\cup L_{m}$ be a link in $S^{3}$ of $m$ components. The boundary of $S^{3}\setminus\nu L$ is orientation-preservingly diffeomorphic to a disjoint union of $m$ tori $\bigcup_{j=1}^{m}S^{1}\times S^{1}$. The diffeomorphism type of the $3$-manifold obtained by Dehn surgery along $L$, i.e. by gluing $m$ solid tori $\bigcup_{j=1}^{m}S^{1}\times D^{2}$ to $S^{3}\setminus\nu L$, is determined by the image of $\bigcup_{j=1}^{m}\{1\}\times\partial D^{2}$ under an orientation-preserving diffeomorphism
@@
is called the surgery coefficient of the link component $L_{j}$. By the above discussion, the diffeomorphism-type of the $3$-manifold obtained by Dehn surgery on the link $L$ is determined by a diagram of $L$ where each component $L_{j}$ of $L$ is labelled by $p_{j}/q_{j}$. We call $L$ together with the labellings $p_{j}/q_{j}$, $i\in\{1,\dots,m\}$, a surgery diagram.
@@
Let us now turn to Rolfsen twists. To that end, let $L=L_{1}\cup\dots\cup L_{m}$ be a surgery diagram with an unknotted component, say $L_{1}$, with surgery coefficient $p_{1}/q_{1}$. For any integer $N\in\mathbb{Z}$, an $N$-fold Rolfsen twist along $L_{1}$ creates out of $L$ a new surgery diagram $L^{\prime}$ as follows:
  • we cut the exterior of $L_{1}$, which is a solid torus, along a Seifert disk of $L_{1}$ and re-glue it with an $N$-fold twist (where $N=1$ corresponds to one right-handed full twist) to get a new link $L^{\prime}=L^{\prime}_{1}\cup\dots\cup L^{\prime}_{m}$ (whose components are in natural bijection with the components of $L$),
@@
If $L$ and $L^{\prime}$ are two surgery diagrams that differ by a Rolfsen twist, then the $3$-manifolds obtained by Dehn surgery along $L$ and $L^{\prime}$, respectively, are orientation-preservingly diffeomorphic.
@@
The process of cutting the exterior of $L_{1}$ along a Seifert disk of $L_{1}$ and re-glueing it with an $N$-fold twist describes a diffeomorphism of the exterior of $L$ to the exterior of $L^{\prime}$. It is straightforward to check that this diffeomorphism acts on the boundary tori as
@@ 3. Garside theory
For each braid $\beta$ on $k$ strands, there is an integer $N\geq 0$ such that  $T_{k}^{N}\beta$ is a positive braid on $k$ strands. More concretely, we can choose $N$ to be the number of negative Artin generators in a braid word representative of $\beta$.
@@ 4. Braid positive surgery diagrams
Given a closed, oriented, connected $3$-manifold $M$, let $L$ be a link whose number of components realizes $s(M)$. We can represent $L$ as the closure of a braid $\beta$ on, say, $k$ strands. Figure 4(a) represents a braid diagram $\beta$ of $L$. Inserting a new component with surgery coefficient $\infty=1/0$ and performing an $N$-fold Rolfsen twist, we obtain the diagram shown in Figure 4(b). Figure 4(c) shows the result of an isotopy of the new link diagram performed on the new component. The link $L^{\prime}$ in Figure 4(c) arises as the closure of a braid $\beta^{\prime}$. Note that $\beta^{\prime}$ is a positive braid whenever the braid $\alpha\vcentcolon=T_{k}^{N}\beta$ on $k$ strands is a positive braid. The theorem thus follows from Lemma 4 choosing $N$ sufficiently large. ∎

### 21 Apr 2026 v3

@@ Abstract.
In this short note, we prove that every closed, oriented, connected 33-manifold arises as Dehn surgery along a braid positive link.
@@ 1. Introduction
A classical result by Lickorish and Wallace [LIC62, WAL61] states that every closed, oriented, connected 33-manifold $M$ arises as Dehn surgery along a link $L$ in $S^{3}$. That is, $M$ can be obtained by removing an open tubular neighbourhood $\nu L$ of $L$ from $S^{3}$ and gluing back to each torus boundary component of $S^{3}\setminus\nu L$ a solid torus by an orientation-preserving diffeomorphism of its boundary. Our motivation in this note stems from the following type of question.
@@
Let $\mathcal{L}$ be a set of links in $S^{3}$. Can each closed, oriented, connected 33-manifold be obtained by Dehn surgery along a link $L\in\mathcal{L}$?
Considering the first homology of a 33-manifold obtained by Dehn surgery directly implies that the number of components of links in $\mathcal{L}$ is unbounded if $\mathcal{L}$ gives a positive answer to Question 1. Question 1 was answered affirmatively for the following sets of links:
@@
In contrast to these results, not every closed, oriented, connected 33-manifold can be obtained by Dehn surgery along a torus link since Dehn surgeries along torus links are never hyperbolic [MOS71, KS10]. Moreover, the answer to Question 1 is negative for the set of Montesinos links by [IJ09] and thus also for pretzel links. Question 1 also has a negative answer for 2-bridge links since these have at most two components.
@@
Every closed, oriented, connected 33-manifold $M$ can be obtained by Dehn surgery along a braid positive link $L$ in $S^{3}$. Moreover, $L$ can be chosen to have at most $s(M)+1$ many components.
@@
All 33-manifolds are assumed to be smooth and oriented, and are considered up to orientation-preserving diffeomorphism. For us, a link in a 33-manifold $M$ is a smooth, closed 11-dimensional submanifold of $M$. Surgery coefficients (see Section 2) are measured with respect to the Seifert framing.
@@ Acknowledgements
This project started at the conference on Trisections and related topics, October 2025, CIRM, Marseille, and was finished at the workshop on Knotted surfaces in 44-manifolds, December 2025, University Regensburg. We thank Stefan Friedl for helpful comments on a first draft.
@@ 2. Surgery diagrams and Rolfsen twists
First, let us briefly review the basics of Dehn surgery. For more details, the reader is referred, for example, to [ROL03]. Let $L=L_{1}\cup\dots\cup L_{m}$ be a link in $S^{3}$ of $m$ components. The boundary of $S^{3}\setminus\nu L$ is orientation-preservingly diffeomorphic to a disjoint union of $m$ tori $\bigcup_{j=1}^{m}S^{1}\times S^{1}$. The diffeomorphism type of the 33-manifold obtained by Dehn surgery along $L$, i.e. by gluing $m$ solid tori $\bigcup_{j=1}^{m}S^{1}\times D^{2}$ to $S^{3}\setminus\nu L$, is determined by the image of $\bigcup_{j=1}^{m}\{1\}\times\partial D^{2}$ under an orientation-preserving diffeomorphism
@@
is called the surgery coefficient of the link component $L_{j}$. By the above discussion, the diffeomorphism-type of the 33-manifold obtained by Dehn surgery on the link $L$ is determined by a diagram of $L$ where each component $L_{j}$ of $L$ is labelled by $p_{j}/q_{j}$. We call $L$ together with the labellings $p_{j}/q_{j}$, $i\in\{1,\dots,m\}$, a surgery diagram.
@@
Let us now turn to Rolfsen twists. To that end, let $L=L_{1}\cup\dots\cup L_{m}$ be a surgery diagram with an unknotted component, say $L_{1}$, with surgery coefficient $p_{1}/q_{1}$. For any integer $N\in\mathbb{Z}$, an NN-fold Rolfsen twist along $L_{1}$ creates out of $L$ a new surgery diagram $L^{\prime}$ as follows:
  • we cut the exterior of $L_{1}$, which is a solid torus, along a Seifert disk of $L_{1}$ and re-glue it with an NN-fold twist (where $N=1$ corresponds to one right-handed full twist) to get a new link $L^{\prime}=L^{\prime}_{1}\cup\dots\cup L^{\prime}_{m}$ (whose components are in natural bijection with the components of $L$),
@@
If $L$ and $L^{\prime}$ are two surgery diagrams that differ by a Rolfsen twist, then the 33-manifolds obtained by Dehn surgery along $L$ and $L^{\prime}$, respectively, are orientation-preservingly diffeomorphic.
@@
The process of cutting the exterior of $L_{1}$ along a Seifert disk of $L_{1}$ and re-glueing it with an NN-fold twist describes a diffeomorphism of the exterior of $L$ to the exterior of $L^{\prime}$. It is straightforward to check that this diffeomorphism acts on the boundary tori as
@@ 3. Garside theory
For each braid $\beta$ on $k$ strands, there is an integer $N\geq 0$ such that  $T_{k}^{N}\beta$ is a positive braid on $k$ strands. More concretely, we can choose NN to be the number of negative Artin generators in a braid word representative of $\beta$.
@@ 4. Braid positive surgery diagrams
Given a closed, oriented, connected 33-manifold $M$, let $L$ be a link whose number of components realizes $s(M)$. We can represent $L$ as the closure of a braid $\beta$ on, say, $k$ strands. Figure 4(a) represents a braid diagram $\beta$ of $L$. Inserting a new component with surgery coefficient $\infty=1/0$ and performing an NN-fold Rolfsen twist, we obtain the diagram shown in Figure 4(b). Figure 4(c) shows the result of an isotopy of the new link diagram performed on the new component. The link $L^{\prime}$ in Figure 4(c) arises as the closure of a braid $\beta^{\prime}$. Note that $\beta^{\prime}$ is a positive braid whenever the braid $\alpha\vcentcolon=T_{k}^{N}\beta$ on $k$ strands is a positive braid. The theorem thus follows from Lemma 4 choosing NN sufficiently large. ∎